# Minimal Ranks, Maximum Confidence: Parameter-efficient Uncertainty Quantification for LoRA

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) enables parameter-efficient fine-tuning of large language models by decomposing weight updates into low-rank matrices, significantly reducing storage and computational overhead. While effective, standard LoRA lacks mechanisms for uncertainty quantification, leading to overconfident and poorly calibrated models. Bayesian variants of LoRA address this limitation, but at the cost of a significantly increased number of trainable parameters, partially offsetting the original efficiency gains. Additionally, these models are harder to train and may suffer from unstable convergence. In this work, we propose a novel parameter-efficient Bayesian LoRA, demonstrating that effective uncertainty quantification can be achieved in very low-dimensional parameter spaces. The proposed method achieves strong performance with improved calibration and generalization while maintaining computational efficiency. Our empirical findings show that, with the appropriate projection of the weight space: (1) uncertainty can be effectively modeled in a low-dimensional space, and (2) weight covariances exhibit low ranks.
@@
Patryk Marszałek  Klaudia Bałazy  Jacek Tabor  Tomasz Kuśmierczyk Jagiellonian University
11footnotetext: Primary contributors.
@@ 1 Introduction
Method $r$ $k$ Parameters
Standard LoRA 2 - 0.2M
@@
Bayesian SWAG-LoRA 2 10 2.4M
@@
Bayesian treatment is then frequently proposed to address overconfidence in neural networks (Blundell et al., 2015; Kristiadi et al., 2020; Aitchison et al., 2021; Izmailov et al., 2021). Consequently, recently proposed Bayesian variants of LoRA (Onal et al., 2024; Robeyns, 2024; Doan et al., 2025) address the aforementioned challenges by introducing uncertainty estimation directly into the fine-tuning process. During training, these models continuously adjust both the mean and covariance of fine-tuned parameters to achieve better generalization and uncertainty quantification.
@@
In the Appendix, we supplement the results presented in the paper with a detailed overview of the experimental setup. We make our code publicly available at github.
@@
Figure 3: Median±$\pm$±std. accuracy (left), ECE (middle), and NLL (right) on 4 GLUE tasks (rows) vs. total parameter count for several methods and varying ranks $r$. B-LoRA-XS (our) achieves the accuracy and the calibration of SWAG-LoRA while using significantly fewer parameters than LoRA. See Fig. 1 for averaged results.
2 Method
LoRA fine-tunes large pre-trained models by learning low-rank weight updates $\Delta W$ instead of training the weights $W$ directly. For a pre-trained parameter matrix $W^{0}\in\mathbb{R}^{m\times n}$ that is kept fixed, LoRA learns a rank-$r$ update $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$ have far fewer parameters. The effective weight is then: $W=W^{0}+\Delta W=W^{0}+AB,$ where only $A$ and $B$ are trained. Typically, LoRA is applied jointly at multiple layers, yielding a set of updates $\{\Delta W_{l}\}$.
@@ 2 Method
Bayesian LoRAs obtain the posterior for $\{\Delta W_{l}\}$ through the learned posterior for $\theta=\cup_{l}\{A_{l}\cup B_{l}\}$, where $l$ indexes the weight updates (layers). The posterior itself is approximated either using a set of particles or a closed-form distribution. Due to its superior performance, we rely on the latter and assume $p(\theta|\mathcal{D})\approx\mathcal{N}(\mu,\Sigma)$, where $\mu$ is the vector of means (of size equal to the number of learned parameters) and $\Sigma$ is the covariance matrix, whose size grows quadratically with the total number of parameters. Notably, we aim to model cross-layer interdependencies, requiring covariance estimation also across weights in different layers $\{l\}$. However, this results in an impractically large number of parameters. Consequently, we explore methods to reduce this cost by representing distributions $p(\{\Delta W\}|\mathcal{D})$ differently.
@@
Figure 4: Impact of the posterior covariance matrix rank ($k=0$ indicates the case with no off-diagonal elements) for CoLA (top) and MRPC (bottom). For brevity, confidence bars ($\pm$ standard deviation) are omitted for MRPC. The colored lines represent non-Bayesian baselines (e.g., standard LoRA or LoRA-XS at a given rank $r$).
B-LoRA-XS, proposed in this paper, leverages the frozen projections $A$ and $B$ for effective and efficient Bayesian learning. In Sec. 3, we empirically demonstrate that $A$ and $B$, obtained from the SVD of the pre-trained weights, are not only effective for point-wise fine-tuning but also enable effective uncertainty quantification for $\{\Delta W_{l}\}$ through modeling covariances for $\{R_{l}\}$. Although we never compute it explicitly, the covariance matrix for individual $\Delta W$ is expressed as $\Sigma_{\Delta W}=(B^{T}\otimes A)\Sigma_{R}(B^{T}\otimes A)^{T}$, where $\Sigma_{R}$ is the (intra-layer) covariance matrix for $R$ and $\otimes$ denotes the Kronecker product.
In practice, we simply learn the joint posterior $p(\theta=\cup_{l}R_{l}|\mathcal{D})\approx\mathcal{N}(\mu,\Sigma)$ for the inner matrices $R$. The covariance matrix $\Sigma$ captures both inter-layer and intra-layer dependencies, allowing the model to learn complex relationships. At inference, similar to LoRA, we use samples of $R$ along with the respective projections $A$ and $B$ to obtain $h$, as realized through samples of $\Delta W$, however without ever computing it explicitly.
We illustrate B-LoRA-XS method in Fig. 2. Our method uses the total of $|\theta|\cdot(k+2)$ parameters, where $|\theta|=\sum_{l}r^{2}_{l}$.
@@ 3 Experiments
Setup: We performed experimental evaluation on four GLUE tasks Wang et al. (2019) (RTE, MRPC, CoLA, and SST-2) using RoBERTA-large Liu et al. (2019). We compare our method (B-LoRA-XS) against the standard LoRA, LoRA-XS – a parameter efficient variant, and against SWAG-LoRA Onal et al. (2024) – a Bayesian variant. For LoRA-XS and B-LoRA-XS we considered ranks $r\in\{2,8,16,25\}$ and for LoRA and SWAG-LoRA due to limited budget we were able to test $r\in\{2,8\}$. The number of parameters (a proxy for storage and computation costs) as a function of ranks $r$ and $k$ is summarized in Fig. 2. We report accuracy (higher is better), ECE and NLL (lower is better) of median±$\pm$±std across 5 runs.
Performance analysis: Fig. 3 compares accuracy and calibration against total parameter count across 4 datasets. Bayesian variants, including B-LoRA-XS (ours) and SWAG-LoRA, outperform their non-Bayesian counterparts, particularly in ECE and NLL. However, our model achieves the best results with 5–15 times fewer parameters.
Apart from very tight budgets (e.g., #parameters $<300k$) where LoRA-XS excels, B-LoRA-XS is also the best option in terms of accuracy. While SWAG-LoRA sometimes performs well, its results vary significantly between runs. In contrast, B-LoRA-XS exhibits stable and consistent convergence. Finally, as results for MRPC and CoLA suggest, its performance remains robust across different values of $k$, whereas SWAG-LoRA’s ECE deteriorates significantly at $k=2$.
@@
4 Related Work
PEFT:
Bayesian LoRAs:
Several Bayesian LoRA methods have been proposed, each employing different Bayesian techniques to address these challenges. SWAG-LoRA Onal et al. (2024) combines Stochastic Weight Averaging-Gaussian (SWAG) with LoRA to enable approximate Bayesian inference, significantly improving model calibration and reducing overconfidence. Laplace-LoRA Robeyns (2024) applies a Laplace approximation to the posterior over LoRA parameters. Bella Doan et al. (2025) introduces an approach that reduces the cost of Bayesian deep ensembles by applying multiple low-rank perturbations to a pre-trained model. BLoB (Bayesian Low-Rank Adaptation by Backpropagation) Wang et al. (2024) jointly learns both the mean and covariance of model parameters throughout the fine-tuning process using Variational Inference. B-LoRA Meo et al. (2024) introduces a Bayesian perspective to both quantization and rank selection by using a prior distribution over these hyperparameters, optimizing model efficiency and reducing bit operations.
@@ 5 Conclusion
Our method consistently achieves lower expected calibration error and negative log-likelihood compared to the standard and parameter-efficient LoRA. Moreover, it matches or surpasses the accuracy of the Bayesian LoRA baseline while using significantly fewer parameters, exhibiting greater training stability, and relying on simpler, lower-rank covariance representations.
We conclude that, with an appropriate weight space transformation, combining low-rank parameter updates with suitably low-rank covariance approximations consistently improves both predictive performance and calibration.
@@ Acknowledgments
Appendix A Scientific Artifacts Licenses
@@
Scientific Artifacts: RoBERTa-large (MIT), MRPC (Unknown), RTE (Unknown), COLA (Unknown), SST-2 (Unknown), HuggingFace Transformers Library (Apache-2.0), SWAG-LoRa repository***https://github.com/fortuinlab/swag-lora (MIT), LoRa-XS repositoryhttps://github.com/MohammadrezaBanaei/LoRA-XS (Unknown).
Appendix B Model Size And Budget
@@
Appendix C Statistics For Data
@@
Appendix D Experimental Setup Details
The study was conducted on a subset of the GLUE benchmark Wang et al. (2019), specifically on RTE, MRPC, CoLA, and SST-2 tasks (with the original train-validation split), using RoBERTa-large Liu et al. (2019) checkpoints from the HuggingFace Transformers library Wolf et al. (2020). For the RTE and MRPC tasks, we followed LoRA-XS and initialized LoRA-XS modules with weights fine-tuned on the MNLI task. We integrated B-LoRA-XS/LoRA-XS modules into the Query, Value, Attention Output, and Output Fully Connected weight matrices in all transformer layers Vaswani et al. (2017), whereas due to budget limits, standard LoRA and SWAG-LoRA modules were added only to the Query and Value matrices. Note, this is sufficient for SWAG-LoRA to achieve its best performance.

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) enables parameter-efficient fine-tuning of large language models by decomposing weight updates into low-rank matrices, significantly reducing storage and computational overhead. While effective, standard LoRA lacks mechanisms for uncertainty quantification, leading to overconfident and poorly calibrated models. Bayesian variants of LoRA address this limitation, but at the cost of a significantly increased number of trainable parameters, partially offsetting the original efficiency gains. Additionally, these models are harder to train and may suffer from unstable convergence. In this work, we propose a novel parameter-efficient Bayesian LoRA via subspace inference, demonstrating that effective uncertainty quantification can be achieved in very low-dimensional parameter spaces. The proposed method achieves strong performance with improved calibration and generalization while maintaining computational efficiency. Our empirical findings show that, with the appropriate projection of the weight space: (1) uncertainty can be effectively modeled in a low-dimensional space, and (2) weight covariances exhibit low ranks.
@@
Patryk Marszałekthanks: Denotes primary contributors.  Klaudia Bałazy  Jacek Tabor  Tomasz Kuśmierczyk11footnotemark: 1thanks: Correspondence: t.kusmierczyk@uj.edu.pl Jagiellonian University
@@ 1 Introduction
Method $r$ $k$ # Parameters
Standard
LoRA 2 - 0.2M
@@
Bayesian
SWAG-LoRA 2 10 2.4M
@@
Bayesian treatment is then frequently proposed to address overconfidence in neural networks (Blundell et al., 2015; Kristiadi et al., 2020; Aitchison et al., 2021; Izmailov et al., 2021). Consequently, recently proposed Bayesian variants of LoRA (Onal et al., 2024; Yang et al., 2024; Doan et al., 2025) address the aforementioned challenges by introducing uncertainty estimation directly into the fine-tuning process. During training, these models continuously adjust both the mean and covariance of fine-tuned parameters to achieve better generalization and uncertainty quantification.
@@
A key contribution of our work lies in introducing Bayesian learning within a low-rank projected subspace derived from pre-trained weights. While the individual components of our method, such as SVD projections and Bayesian inference, are established techniques, their synergistic application to learn Bayesian posteriors within a compressed subspace constitutes a meaningful conceptual innovation. Our approach enables uncertainty-aware fine-tuning with strong parameter efficiency, yielding improved uncertainty quantification with minimal computational overhead.
In the Appendix, we supplement the results presented in the paper with a discussion of related work, a detailed overview of the experimental setup, and the exact numeric values for the figures in the main text.
The source code is available online [^1] Source code: https://github.com/gmum/b-lora-xs.
@@
2 Method: B-LoRA-XS
LoRA fine-tunes large pre-trained models by learning low-rank weight updates $\Delta W$ instead of training the weights $W$ directly. For a pre-trained parameter matrix $W^{0}\in\mathbb{R}^{m\times n}$ that is kept fixed, LoRA learns a rank-$r$ update $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$ have far fewer parameters. The effective weight is then: $W=W^{0}+\Delta W=W^{0}+AB,$ where only $A$ and $B$ are trained. LoRA is then typically applied jointly for multiple layers $l$, yielding a set of updates $\{\Delta W_{l}\}$. Then, Bayesian treatment of LoRA can improve its calibration and uncertainty quantification.
Bayesian treatment of a neural network involves finding the posterior $p(\theta\mid\mathcal{D})$ given training data $\mathcal{D}$. By Bayes’ theorem: $p(\theta\mid\mathcal{D})\;=\;\frac{p(\mathcal{D}\mid\theta)\,p(\theta)}{p(\mathcal{D})},$ where $\theta$ represents the model’s parameters (i.e., weights) considered random variables. Specifically, for the Bayesian LoRA setting, $\theta$ denotes a set of the learned model updates, while the remaining frozen weights are hidden inside the model likelihood, given by $p(\mathcal{D}\mid\theta)=\prod_{i\in[\mathcal{D}]}p(y_{i}|x_{i},\theta)$. The learned posterior allows Bayesian model averaging at inference as: $p(y_{*}\mid x_{*},\mathcal{D})\;=\;\int p(y_{*}\mid x_{*},\theta)\,p(\theta\mid\mathcal{D})\,d\theta\approx\frac{1}{S}\sum_{\theta\sim p(\theta|D)}p(y_{*}\mid x_{*},\theta).$
Bayesian LoRAs obtain the posterior for $\{\Delta W_{l}\}$ through the learned posterior for $\theta=\cup_{l}\{A_{l}\cup B_{l}\}$, where $l$ indexes the weight updates (layers). The posterior itself is approximated either using a set of particles or a closed-form distribution. Due to its superior performance, we rely on the latter and assume $p(\theta|\mathcal{D})\approx\mathcal{N}(\mu,\Sigma)$, where $\mu$ is the vector of means (of size equal to the number of learned parameters) and $\Sigma$ is the covariance matrix, whose size grows quadratically with the total number of parameters. Notably, we aim to model cross-layer interdependencies, requiring covariance estimation also across weights in different layers $\{l\}$. This however results in an impractically large number of parameters. Consequently, we explore methods to reduce this cost by representing distributions $p(\{\Delta W_{l}\}|\mathcal{D})$ differently, e.g., using SVD-based projections.
@@ 2 Method: B-LoRA-XS
B-LoRA-XS, proposed in this paper, leverages the frozen projections $A$ and $B$ for effective and efficient Bayesian learning. Its core idea is to apply the Bayesian treatment in the extremely compressed parameter space given by SVD-based projections, making Bayesian inference tractable and highly efficient. Although this specific idea is novel, it can be related to the framework of subspace Bayesian Inference.
Subspace Inference (SI) (Izmailov et al., 2020) was proposed as a remedy for the intractability of full-dimensional posteriors in modern networks. Low-dimensional affine manifold, carefully centred on a well-trained solution, already contains a rich family of high-performing weight vectors. Performing Bayesian integration restricted to that manifold restores calibrated uncertainty without revisiting the entire parameter space. In particular, given a well-trained reference point $\bar{w}\in\mathbb{R}^{d}$ and $K\ll d$ orthonormal basis vectors-e.g. the leading PCA directions of an SGD trajectory - the learning subspace is $\mathcal{S}:=\bigl{\{}w=\bar{w}+Pz\bigm{|}z\in\mathbb{R}^{K}\bigr{\}}$, $P\in\mathbb{R}^{d\times K},$ where Bayesian parameters are the low-dimensional representations $\theta=\{z\}$.
B-LoRA-XS defines a layer-local manifold that leaves the pretrained backbone untouched. For each frozen weight matrix $W^{0}_{l}$, an SVD $W^{0}_{l}=U_{l}S_{l}V_{l}^{T}$ is computed once, and the top $r$ singular directions yield fixed projectors $A_{l}=U_{l,r}S_{l,r}$ and $B_{l}=V_{l,r}^{T}$. Note that, whereas SI learns $P$ from the training dynamics, B-LoRA-XS projects onto directions already favored by the pretrained backbone. Then, all task-specific variability is captured by these square adapters $R_{l}\in\mathbb{R}^{r\times r}$.
Vectorising and stacking these layer-projections defines $w=\bar{w}+P_{B}z_{B}$, $P_{B}=\mathrm{blockdiag}\bigl{(}B_{l}^{T}\otimes A_{l}\bigr{)},$ where $w=[\mathrm{vec}(W_{1})^{T},\ldots,\mathrm{vec}(W_{l})^{T},\ldots]^{T}$, $\bar{w}=[\mathrm{vec}(W_{1}^{0})^{T},\ldots,\mathrm{vec}(W_{l}^{0})^{T},\ldots]^{T}$, and $P_{B}$ is a block-diagonal matrix with blocks $B_{l}^{T}\otimes A_{l}$, one for each layer $l$.
Thus B-LoRA-XS explores an affine subspace $\mathcal{S}_{B}:=\{w=w^{0}+P_{B}z_{B}|z_{B}\in\mathbb{R}^{\sum_{l}r^{2}}\},$ whose dimension scales with $r^{2}$ per adapted layer.
In Sec. 3, we empirically demonstrate that $A_{l}$ and $B_{l}$, obtained from the SVD of the pre-trained weights, are not only effective for point-wise fine-tuning but also enable effective uncertainty quantification for $\{\Delta W_{l}\}$ through modeling covariances for $\{R_{l}\}$. Although we never compute it explicitly, the covariance matrix for individual $\Delta W$ is expressed as $\Sigma_{\Delta W_{l}}=(B_{l}^{T}\otimes A_{l})\Sigma_{R_{l}}(B_{l}^{T}\otimes A_{l})^{T},$, where $\Sigma_{R}$ is the (intra-layer) covariance matrix for $R$ and $\otimes$ denotes the Kronecker product.
In practice, we simply learn the joint posterior $p(\theta=\cup_{l}R_{l}|\mathcal{D})\approx\mathcal{N}(\mu,\Sigma)$ (only) for the inner matrices $R$. The covariance matrix $\Sigma$ captures both inter-layer and intra-layer dependencies, allowing the model to learn complex relationships. At inference, similar to LoRA, we use samples of $R$ along with the respective projections $A$ and $B$ to obtain $h$, as realized through samples of $\Delta W$, however without ever computing it explicitly.
We illustrate B-LoRA-XS method in Fig. 2. Our method uses the total of $|\theta|\cdot(k+2)$ parameters, where $|\theta|=\sum_{l}r^{2}$.
@@
Figure 4: Impact of the posterior covariance matrix rank ($k=0$ indicates the case with no off-diagonal elements) for CoLA (top) and MRPC (bottom). For brevity, confidence bars ($\pm$ standard deviation) are omitted for MRPC. The colored lines represent non-Bayesian baselines (e.g., standard LoRA or LoRA-XS at a given rank $r$). The exact numeric values underlying the plots are reported in Tables 5 and 6 in the Appendix.
@@ 3 Experiments
[…]
Performance analysis: Fig. 3 compares accuracy, Expected Calibration Error (ECE), and Negative Log-Likelihood (NLL) against total parameter count across 4 datasets. Our main claim is that B-LoRA-XS improves overall model performance, with a particular focus on calibration metrics. Indeed, Figure 3 (middle and right) demonstrates that B-LoRA-XS consistently yields lower ECE and NLL compared to standard LoRA across all parameter scales. Regarding accuracy (Figure 3: left), while standard LoRA shows marginally better results for a few configurations at moderate parameter scales, the majority of configurations show B-LoRA-XS matching or exceeding the accuracy of standard LoRA. More importantly, in no setting does standard LoRA significantly outperform B-LoRA-XS in terms of calibration, which is a primary focus of our work. Bayesian variants, including B-LoRA-XS and SWAG-LoRA, generally outperform their non-Bayesian counterparts in ECE and NLL. However, our model achieves these strong calibration results with 5–15 times fewer parameters than SWAG-LoRA. Moreover, while SWAG-LoRA sometimes performs well, its results vary significantly between runs. In contrast, B-LoRA-XS exhibits stable and consistent convergence. Finally, as results for MRPC and CoLA suggest, its performance remains robust across different values of $k$, whereas SWAG-LoRA’s ECE deteriorates significantly at $k=2$.
@@ 4 Conclusion
Our method’s primary strength lies in its calibration capabilities; it consistently achieves lower expected calibration error and negative log-likelihood compared to standard LoRA and LoRA-XS across various parameter scales. While standard LoRA may exhibit marginally better accuracy in a few specific configurations, B-LoRA-XS generally matches or exceeds its accuracy in most settings, and critically, always provides superior or equal calibration. Compared to the Bayesian LoRA baseline, B-LoRA-XS matches or surpasses its accuracy and calibration performance while using significantly fewer parameters, exhibiting greater training stability, and relying on simpler, lower-rank covariance representations.
Limitations
While B-LoRA-XS demonstrates promising results in parameter-efficient uncertainty quantification, several limitations should be acknowledged. First, the effectiveness of B-LoRA-XS inherently depends on the quality of the initial SVD projection derived from pre-trained weights (as in LoRA-XS). If the principal components of the pre-trained model are not well-aligned with the requirements of a significantly different downstream task, the performance might be suboptimal. Second, our method employs SWAG with a low-rank approximation for the covariance matrix. While efficient, this is one specific approach to approximate Bayesian inference. Other techniques (e.g., more sophisticated variational inference methods or different posterior approximations) might yield different trade-offs between performance, calibration, and computational cost, and were not explored in this work. Third, although B-LoRA-XS significantly reduces the number of trainable parameters for Bayesian adaptation, the inference process still requires multiple forward passes for sampling, which increases computational cost compared to non-Bayesian LoRA or LoRA-XS. Fourth, our empirical validation is conducted on GLUE classification tasks using RoBERTa-Large. The generalizability of B-LoRA-XS’s benefits to other model architectures, much larger model scales, or different task types (such as text generation or more complex reasoning tasks) warrants further investigation. Finally, the optimal choice of LoRA rank $r$ and SWAG covariance rank $k$ might require careful tuning for different datasets and models, potentially adding to the practical overhead of applying the method effectively.
@@ Acknowledgments
Appendix A Related Work
PEFT:
Bayesian LoRAs:
Several Bayesian LoRA methods have been proposed, each employing different Bayesian techniques to address these challenges. SWAG-LoRA Onal et al. (2024) combines Stochastic Weight Averaging-Gaussian (SWAG) with LoRA to enable approximate Bayesian inference, significantly improving model calibration and reducing overconfidence. Laplace-LoRA Yang et al. (2024) applies a Laplace approximation to the posterior over LoRA parameters. Bella Doan et al. (2025) introduces an approach that reduces the cost of Bayesian deep ensembles by applying multiple low-rank perturbations to a pre-trained model. BLoB (Bayesian Low-Rank Adaptation by Backpropagation) Wang et al. (2024) jointly learns both the mean and covariance of model parameters throughout the fine-tuning process using Variational Inference. B-LoRA Meo et al. (2024) introduces a Bayesian perspective to both quantization and rank selection by using a prior distribution over these hyperparameters, optimizing model efficiency and reducing bit operations.
Appendix B Scientific Artifacts Licenses
@@
Scientific Artifacts: RoBERTa-large (MIT), MRPC (Unknown), RTE (Unknown), CoLA (Unknown), SST-2 (Unknown), HuggingFace Transformers Library (Apache-2.0), SWAG-LoRA repository [^2] https://github.com/fortuinlab/swag-lora (MIT), LoRA-XS repository [^3] https://github.com/MohammadrezaBanaei/LoRA-XS (Unknown).
Appendix C Model Size And Budget
@@
Appendix D Statistics For Data
@@
Appendix E Experimental Setup Details
The study was conducted on a subset of the GLUE benchmark Wang et al. (2018), specifically on RTE, MRPC, CoLA, and SST-2 tasks (with the original train-validation split), using RoBERTa-large Liu et al. (2019) checkpoints from the HuggingFace Transformers library Wolf et al. (2020). For the RTE and MRPC tasks, we followed LoRA-XS and initialized LoRA-XS modules with weights fine-tuned on the MNLI task. We integrated B-LoRA-XS/LoRA-XS modules into the Query, Value, Attention Output, and Output Fully Connected weight matrices in all transformer layers Vaswani et al. (2017), whereas due to budget limits, standard LoRA and SWAG-LoRA modules were added only to the Query and Value matrices. Note this is sufficient for SWAG-LoRA to achieve its best performance.